If money can be invested at 7% compounded monthly, find the present value of which is due after 2 years and 11 months from today.
Answer by Theo(13342)

the general formula is:
f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

if the time period is in months, then the interest rate per time period is per month and the number of time periods is months.

to find the present value, do the following:
start with f = p * (1 + r) ^ n
solve for p to get:
p = f / ((1 + r) ^ n)

if the rate per time period is given in percent per year, and the time periods is in months, then do the following:
r% per year = 7
divide that by 100 to get r per year = .07
divide that by 12 to get r per month = .07/12.

if the number of time periods is given in number of years, and the number of time periods needs to be in months, then do the following:
2 years and 11 months is equal to 24 months and 11 months which is equal to 35 months.

you have r = .07/12 per month and n = 35 months.

the formula of p = f / ((1 + r) ^ n) becomes:
p = f / ((1 + .07/12) ^ 35
simplify to get:
p = f / 1.225775232

your present value will be the future value divided by 1.225775232.

for example, if the future value is 50,000, then the present value will be 50,000 / 1.225775232 = 40790.51256